how will you factor 2x^2-178x-14280=0
(2x ) (x ) << I cant find the factors

Answer by stanbon(75887)

how will you factor 2x^2-178x-14280=0
(2x ) (x ) << I cant find the factors
-------------------------------------------
(x-140)(2x+102)
--------------------
How to do these?
Consider b^2-4ac = 178^2-4*2*-14280 = 145924
Take the sqrt to get 382
Add -b = 178 to get 560
Divide by 2a = 4 to get 140
So, one of the factors is (x-140)
Divide 14280 by 140 to get 102
----------
The other factor must be 2x+102
Check it out to see if it is correct.
========================================
Cheers,
Stan H.